b) Find whether the following series are convergent or divergent √ 𝟏 /πŸ’ + √ 𝟐 /πŸ” + √ πŸ‘ /8 + ...


1
Expert's answer
2022-03-15T10:20:34-0400

"\sqrt{\frac{1}{4}}+\sqrt{\frac{2}{6}}+\sqrt{\frac{3}{8}}+...=\sum_{n=1}^{\infty}\sqrt{\frac{n}{2(n+1)}}\\
a_n=\sqrt{\frac{n}{2(n+1)}}, a_{n+1}=\sqrt{\frac{n+1}{2(n+1+1)}}=\sqrt{\frac{n+1}{2(n+2)}}\\
\lim\limits_{n\to\infty}\frac{a_{n+1}}{a_{n}}=\lim\limits_{n\to\infty}\frac{\sqrt{\frac{n+1}{2(n+2)}}}{\sqrt{\frac{n}{2(n+1)}}}=\\
=\lim\limits_{n\to\infty}\frac{n+1}{\sqrt{n(n+2)}}=1\\
\lim\limits_{n\to\infty}a_n=\lim\limits_{n\to\infty}\sqrt{\frac{n}{2(n+1)}}=\\
=\lim\limits_{n\to\infty}\sqrt{\frac{n}{2n(1+\frac{1}{n})}}=\sqrt{\frac{1}{2}}\neq0"

The series are divergent.